OpenStudy (anonymous):

Determine the maximum value of the function f(x)=4e^(-x) - e^x on the interval -1<=x<=2. I first found the derivative: f'(x)=-4e^(-x) - e^x Then I set it two zero. My problem is that I don't know how to isolate x

OpenStudy (anonymous):

-4e^-x -e^x = 0 e^x = -4 e^-x e^x/e^-x = -4 e^2x = -4

OpenStudy (anonymous):

-4e^(-x) - e^(x) = 0 e^(x) = -4e^(-x) e^(2x) = -4 2x = ln(-4) x = 0.5*ln(-4), not real number. That means that there is no maximum or minimum. BUT, we have an interval [-1, 2], so we just need to check the maximum in the boundaries. f(-1) = 10.505... f(2) = -6.8477... So, the maximum for the given interval is when x = -1

OpenStudy (anonymous):

Oh, i got it now thanks!

OpenStudy (anonymous):

No problem ;)

OpenStudy (anonymous):

I have a question. Was -1 and 2 subbed into the derivative?

OpenStudy (anonymous):

No, the derivative is used to find if there is a point in which the graph as slope 0 and that doesn't happen with this function

OpenStudy (anonymous):

Oh ok i fixed my mistake!

OpenStudy (anonymous):

so, we have to test the boundaries of f(x) and check the local max or min in a interval

OpenStudy (anonymous):

Ok thanks again!